OpenStudy (anonymous):

The legs of an isosceles triangle have lengths x+1 and -x+7. The base has length 3x-3. What is the length of the base? a.6 b.6 c.3 d. cannot be determined i need help i dont even know what to do

OpenStudy (anonymous):

i think the answer is c not sure

OpenStudy (mvpriest):

y do u think its C.

OpenStudy (anonymous):

because (3)+1=4 and (-3)+7= 4 am i wrong?

OpenStudy (mvpriest):

good You know the concept and yes it is C.

OpenStudy (anonymous):

thanks.

OpenStudy (mvpriest):

no problem any time

OpenStudy (freckles):

well it says the length of the base

OpenStudy (freckles):

it is not asking for x

OpenStudy (freckles):

but you did find x

OpenStudy (freckles):

so that is a step in the right direction

OpenStudy (anonymous):

so c is not the answer?

OpenStudy (freckles):

is x the length of the base or is 3x-3 the length of the base?

OpenStudy (mvpriest):

3x-3 is

OpenStudy (freckles):

so you already determined x

OpenStudy (freckles):

now to find the length of the base plug that into 3x-3

OpenStudy (anonymous):

thats 9?

OpenStudy (mvpriest):

okay then oh yea I see

OpenStudy (anonymous):

Equate x+1 to -x+7 and then solve the result for x. x = 3 Replace x in the base length expression with 3 and evaluate the result, which is 6. The answer is a. or b.

OpenStudy (freckles):

3(3)-3=?

OpenStudy (mvpriest):

but yea thats what i got now so... its D.?

OpenStudy (freckles):

9-3 isn't 9

OpenStudy (anonymous):

that is 6

OpenStudy (freckles):

3(3)-3 9-3 <--since 3(3)=9 6

OpenStudy (mvpriest):

wait you put me off track what?

OpenStudy (anonymous):

where did you get the three

OpenStudy (freckles):

you got the 3 remember?

OpenStudy (freckles):

you said x=3

OpenStudy (mvpriest):

oh right ok im back on

OpenStudy (anonymous):

3x-3 the x was 4 from the other solutions? whwere are you getting the three?

OpenStudy (freckles):

 -x+7=x+1 add x on both sides 7=2x+1 subtract 1 on both sides 6=2x x=3 so the bottom (the base) the 3x-3 is 9 since 3(3)-3=9-3=6